OpenStudy (anonymous):

if secA=17/8, verify that (3-4sin^2A)/(4cos^2A-3)=(3-tan^2A/1-3tan^2A)

OpenStudy (shinalcantara):

\[\sec A =\frac{ 17 }{ 8 } is equal \to \frac{ 1 }{ \cos A } = \frac{ 8 }{ 17 }\]

OpenStudy (shinalcantara):

Now try to recall SOH CAH TOA and identify all the values for the 3 sides of a right triangle

OpenStudy (shinalcantara):

i somehow interchanged the numerator and denominator..

OpenStudy (shinalcantara):

\[\frac{ 1 }{ \cos A } = \frac{ 17 }{ 8 } \]

OpenStudy (shinalcantara):